Find the standard form of each equation by completing the square. x^2-2x-20y-79=0
Answer by stanbon(75887)

Find the standard form of each equation by completing the square.
x^2-2x-20y-79 = 0
-----
x^2-2x+1 = 20y+79+1
(x-1)^2 = 20y+80
----
(x-1)^2 = 20(y+4)
----
This is the vertex form for the parabola.
===============
Cheers,
Stan H.